# Towards Scalable and Stable Parallelization of Nonlinear RNNs

### v1

@@ 6.1 Quasi-DEER for Evaluation
We replicate the first experiment from Lim et al. (2024) — evaluating an untrained GRU — and show the results in Figure 2. The experimental task is to evaluate an untrained GRU across a range of hidden state sizes ($D$) and sequence lengths ($T$) on a 16 GB V100 GPU; the inputs to the RNN also have dimension $D$. We compare the wall-clock time and memory usage of three methods: sequential evaluation, DEER, and quasi-DEER.
@@
We show the results in Figure 2. Both DEER and quasi-DEER are as much as twenty times faster than sequential evaluation. We also see that quasi-DEER requires as much as an order of magnitude less memory than DEER, thus allowing the extension of the method into regimes previously infeasible for DEER. Note that the runtimes are similar between DEER and quasi-DEER for small networks, because although quasi-DEER steps are faster, DEER takes fewer iterations to converge. For larger networks, the difference in runtime is more pronounced. In Figure 6 of Appendix B.1.1 we show that in smaller $T$ and $D$ regimes we observe the expected sublinear time scaling with sequence length. This experiment confirms that quasi-DEER can replicate the performance of DEER, but with a smaller memory footprint.
@@ 6.2 Quasi-DEER for Training
We verify that quasi-DEER expedites training nonlinear RNN models. We replicate the third experiment from Lim et al. (2024), where a GRU is trained to classify C. elegans specimens from the time series of principal components of the worms’ body posture (Brown et al., 2013).
We show results in Figure 3. We see that the training dynamics under quasi-DEER leads to the similar validation accuracy trajectories, indicating there is no significant difference in training. However, every quasi-DEER training step is faster by a factor of $2.5$, despite performing $2$-$2.5$ times more Newton iterations per training step. This highlights how quasi-DEER can be used as a drop-in replacement for DEER when training nonlinear RNNs, bringing both time and memory savings.
DEER is prone to “spikes”, where orders of magnitude more steps are required for convergence (Figure 3, Middle Panel). While quasi-DEER is not as susceptible to these spikes (never more than half an order of magnitude), these instabilities motivate the study of stabilizing methods.
Figure 3: Training a GRU with DEER. Comparison of DEER and quasi-DEER during GRU training for the C. elegans time-series classification task (Section 6.2). Each time series has length $T=17,984$. We show the median, and 5-95% interval across a rolling window of 20 training steps. (Left) DEER and quasi-DEER have the similar validation accuracy trajectories, indicating similar training dynamics. The sequential trace shown is for 24 hours of training (compared to 11 and 4 for the whole DEER and quasi-DEER traces). (Center) Each quasi training iteration is 2.5 times faster than each DEER training iteration. Sequential training steps took more than 6 seconds each (not pictured). (Right) Each quasi training iteration requires (approximately) 2-2.5 times more Newton iterations to converge, indicating that each quasi Newton step is approximately 6 times faster than the corresponding DEER Newton step.
@@ 6.3 ELK and Quasi-ELK for Evaluating Autoregressive RNNs
We conclude by studying an application where these numerical instabilities in DEER are critical. We use a small autoregressive GRU (hidden dimension $N_{h}=3$), where the previous sampled value is input into the GRU at the next step. Such autoregressive architectures were not examined by Lim et al. (2024), but are an important class of models. We describe the precise details of the AR GRU we use in Appendix B.3. Crucially, the effective state used by all four methods must be expanded to include the current sampled output value, as well as the current GRU state.
@@
We first repeat the analysis in Section 6.1 (and similar to the evaluation in Lim et al. (2024)) for evaluating a randomly initialized autoregressive GRU. We see in the top left panel of Figure 4 that all four parallelized methods converge rapidly and stably to the correct trace, indicated by a low mean absolute discrepancy (MAD) between the true trace and the generated trace.
@@
We then study a pre-trained GRU that generates a noisy sine wave (see Figure 4, Bottom Panel). The linear recurrence relation eq. 6 was numerically unstable in DEER and Quasi-DEER. To remedy these instabilities in DEER and quasi-DEER, we take the approach described earlier of setting the unstable parts of the trace to a fixed value (here zero). As described earlier, this reliably provides convergence; but at the cost of “resetting” the optimization for large swathes of the trace (Figure 4, Bottom Panel), slowing convergence (see Figure 4, Top Right Panel). This finding highlights how the instabilities of DEER — which are inherited from both pathologies of Newton’s method and the parallel recurrence — can be crippling in even very simple scenarios. While the resetting heuristic does facilitate convergence, the resulting convergence is drastically slower.
Figure 4: ELK stabilizes parallel evaluation of an AR GRU. (Top Left) The mean absolute difference (MAD) evaluated on the outputs converges rapidly for all four methods on a sequence generated by an untrained AR GRU. (Top Right) The MAD for evaluating a trained AR GRU. Undamped DEER variants are unstable and converge slowly. Damping via ELK stabilizes and accelerates convergence. (Bottom) The output after 1, 100, 1000, and 2000 Newton iterations. The black dotted line is the true trace. ELK and quasi-ELK converge rapidly, but DEER and quasi-DEER are unstable. The lines where DEER and quasi-DEER are zero show the heuristic of replacing NaN with zero.
@@
However, on this task, all the parallelized methods (including DEER) have higher wall-clock time than sequential generation. Quasi-ELK is the fastest parallel method, taking 221 milliseconds for convergence, compared to sequential evaluation, taking 96 milliseconds. For comparison, DEER took 1,225 milliseconds. Quasi-ELK therefore still represents a large improvement in runtime over previous methods. We provide timing details and further discussion in Appendix B.3.2.
@@ 7 Conclusion
We proposed methods for scalable and stable parallel evaluation of nonlinear RNNs. DEER (Lim et al., 2024) achieves speedups over sequential evaluation, but incurs quadratic memory, cubic work, and can be numerically unstable. We therefore extended DEER to use quasi-Newton approximations, and provided a novel proof that both DEER and quasi-DEER converge globally. To stabilize DEER, we leveraged a connection between the Levenberg-Marquardt method and Kalman smoothing to enable parallel evaluation of RNNs, allowing us to stabilize DEER while still leveraging fast parallel filtering. We verified empirically that quasi-DEER, ELK, and quasi-ELK improve convergence across a range of metrics and examples. This result allows parallel evaluation of nonlinear RNNs to be scaled beyond what is possible with DEER. Our experiments also highlighted avenues for future research, such as leveraging structured approximations of the residual Jacobian, that still admit fast parallelism but are more accurate approximations, would allow for more accurate steps to be taken in quasi-DEER and quasi-ELK.
When selecting an approach, we offer the following advice: If rapid convergence is reliably observed, our experiments show that quasi-DEER provides the fastest convergence in terms of wall-clock time. If numerical stability is a concern, then ELK offers the most stable performance and convergence in the fewest Newton iterations, but quasi-ELK could be faster in wall-clock time and just as stable.
Acknowledgements
We thank John Duchi and the members of the Linderman Lab for helpful feedback. This work was supported by grants from the NIH BRAIN Initiative (U19NS113201, R01NS131987, & RF1MH133778), the NSF/NIH CRCNS Program (R01NS130789). X.G. would also like to acknowledge support from the Walter Byers Graduate Scholarship from the NCAA. S.W.L. is supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation. The authors have no competing interests to declare.
@@ A.1 Proof of Proposition 1
Proposition 1 If the Jacobians of the dynamics function $f$ are finite everywhere in the state space, then undamped Newton’s method will converge to the true solution, $\mathbf{s}^{*}$, of the fixed point eq. 2 in at most $T$ Newton iterations, for any initial $\mathbf{s}^{(0)}$.
@@
Furthermore, we observe that if $\mathbf{s}_{t}^{(i)}=\mathbf{s}_{t}^{*}$ for all $t$ less than some $t^{(i)}$, then $\mathbf{r}_{t}(\mathbf{s}^{(i)})=\mathbf{0}$ for all $t<t^{(i)}$ by the definition of the residual in eq. 1. Therefore, as long as the Jacobians $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ are finite, we see that the DEER linear recurrence relation necessarily gives $\Delta\mathbf{s}^{(i+1)}_{t}=\mathbf{0}$ for all $t<t^{(i)}$. Furthermore, because $\mathbf{s}_{t^{(i)}}^{*}=f(\mathbf{s}_{t}^{(i)})$, it follows that $\Delta\mathbf{s}_{t^{(i)}}^{(i+1)}=-\mathbf{r}_{t^{(i)}}(\mathbf{s}^{(i)})=\mathbf{s}^{*}_{t^{(i)}}-\mathbf{s}^{(i)}_{t^{(i)}}$. Thus, it follows that after applying another Newton iteration that $s_{t}^{(i+1)}=s_{t}^{*}$ for all $t<t^{(i)}+1$. The global convergence result and bound on Newton iterates now follows by induction. ∎
A.2 The Merit Function has no Local Minima
@@
The merit function $\mathcal{L}(s)$ defined in (7) has a global minimum in the true trace $\mathbf{s}^{*}$, satisfying $\mathcal{L}(s^{*})=0$. It has no other local minima, i.e. no $\mathbf{s}$ such that $\nabla\mathcal{L}(\mathbf{s})=\mathbf{0}$ other than at the unique $\mathbf{s}^{*}$ for which $\mathbf{r}(\mathbf{s}^{*})=\mathbf{0}$.

### v3

@@ 6.1 Quasi-DEER for Evaluation
We first use an experimental design from Lim et al. [1]. The task is to evaluate an untrained GRU across a range of hidden state sizes ($D$) and sequence lengths ($T$) on a 16GB V100 GPU; the inputs to the RNN also have dimension $D$. We compare the wall-clock time and memory usage of three methods: sequential evaluation, DEER, and quasi-DEER. Results are shown Figure 2.
@@
Both DEER and quasi-DEER are up to twenty times faster than sequential evaluation. The runtimes are similar between DEER and quasi-DEER for small networks, because although quasi-DEER steps are faster, quasi-DEER takes more iterations to converge. For larger networks, the difference in runtime is more pronounced. We also see that quasi-DEER requires as much as an order of magnitude less memory than DEER, thus allowing the application to architectural regimes previously infeasible with DEER. In Figure 6 of Appendix B.1.1 we show that in smaller $T$ and $D$ regimes we observe the expected sublinear time scaling with sequence length. This experiment confirms that quasi-DEER can replicate the performance of DEER, but with a smaller memory footprint.
@@ 6.2 Quasi-DEER for Training
We verify that quasi-DEER expedites training nonlinear RNN models. We replicate the third experiment from Lim et al. [1], where a GRU is trained to classify C. elegans phenotypes from the time series of principal components of the worms’ body posture [73].
We show results in Figure 3. We see that the training dynamics under quasi-DEER leads to the similar validation accuracy trajectories. However, every quasi-DEER training step is faster by a factor of $2.5$, despite performing around $2$ times more Newton iterations per training step. This finding highlights how quasi-DEER can replace DEER when training nonlinear RNNs, yielding both time and memory savings. In our experiment, we use the quasi-DEER approximation for the backward pass as well, leading to gradients that are different from DEER in this setting. However, we find that there is negligible degradation in performance (Figure 3, left).
DEER is prone to “spikes”, where orders of magnitude more steps are required for convergence (Figure 3, right). While quasi-DEER is not as susceptible to these spikes (never more than half an order of magnitude), these instabilities motivated the study of stabilizing methods.
Figure 3: Training a GRU with DEER. Comparison of DEER and quasi-DEER during GRU training for the C. elegans time-series classification task (Section 6.2). Each time series has length $T=17,984$. We show the median, and 5-95% interval across a rolling window of 20 training steps. (Left) DEER and quasi-DEER have the similar validation accuracy trajectories, indicating similar training dynamics. The sequential trace shown is for 24 hours of training (compared to 11 and 4 hours for the whole DEER and quasi-DEER traces). (Center) Each quasi training iteration is 2.5 times faster than each DEER training iteration. Sequential training steps took more than 6 seconds each (not pictured). (Right) Each quasi training iteration requires (approximately) 2 times more Newton iterations to converge, indicating that each quasi Newton step is approximately 5 times faster than the corresponding DEER Newton step.
@@ 6.3 ELK and Quasi-ELK for Evaluating Autoregressive RNNs
We conclude by studying an application where these numerical instabilities in DEER are critical. We use a small autoregressive GRU (hidden dimension $N_{h}=3$), where the previous sampled value is input into the GRU at the next step. Such autoregressive architectures were not examined by Lim et al. [1], but are an important class of models. We describe the precise details of the AR GRU we use in Appendix B.3. Crucially, the Markovian state $\mathbf{s}_{t}$ used by all four parallelized methods must be expanded to include the current sampled output value, as well as the current GRU state.
@@
We first repeat the analysis in Section 6.1 (and similar to the evaluation in Lim et al. [1]) for evaluating a randomly initialized autoregressive GRU. We see in the top left panel of Figure 4 that all four parallelized methods converge rapidly and stably to the correct trace, indicated by a low mean absolute discrepancy (MAD) between the true trace and the generated trace.
@@
We then study a pre-trained GRU that generates a noisy sine wave (see Figure 4, bottom). The linear recurrence relation (6) was numerically unstable in DEER and quasi-DEER. To remedy these instabilities, we take the approach described earlier of setting the unstable parts of the trace to a fixed value (here zero). Doing so ensures convergence, but at the cost of “resetting” the optimization for large swathes of the trace (Figure 4, bottom) and slowing convergence (see Figure 4, top right). This finding highlights how the instabilities of DEER — which are inherited from both pathologies of Newton’s method and the parallel recurrence — can be crippling in even very simple scenarios. While resetting allows for convergence, the resulting convergence is very slow.
Figure 4: ELK stabilizes parallel evaluation of an AR GRU. (Top Left) The mean absolute difference (MAD) evaluated on the outputs converges rapidly for all four methods on a sequence generated by an untrained AR GRU. (Top Right) The MAD for evaluating a trained AR GRU. Undamped DEER variants are unstable and converge slowly (using the reset heuristic). ELK stabilizes and accelerates convergence. (Bottom) The output after 1, 100, 1000, and 2000 Newton iterations. The black dotted line is the true trace. ELK and quasi-ELK converge rapidly, but DEER and quasi-DEER are unstable. The lines where DEER and quasi-DEER are zero depict the zeroing heuristic.
@@
However, on this task, all parallelized methods (including DEER) are slower than sequential generation. Quasi-ELK is the fastest parallel method, taking 221 milliseconds, compared to sequential evaluation, taking 96 milliseconds. For comparison, DEER took 1,225 milliseconds. Quasi-ELK therefore still represents a large improvement in runtime over previous parallel methods. We provide timing details and further discussion in Appendix B.3.2.
@@ 7 Conclusion
In this paper we proposed methods for scalable and stable parallel evaluation of nonlinear RNNs. DEER [1] achieved speedups over sequential evaluation, but incurred quadratic memory, cubic work, and numerical instabilities. We therefore extended DEER to use quasi-Newton approximations that reduced computational complexity, and we provided a novel proof that both DEER and quasi-DEER converge globally. To stabilize DEER, we leveraged a connection between the Levenberg-Marquardt method and Kalman smoothing to enable parallel evaluation of RNNs, allowing us to stabilize DEER while still leveraging fast parallel filtering. We empirically verified that quasi-DEER, ELK, and quasi-ELK improve convergence across a range of metrics and examples. This result allows parallel evaluation of nonlinear RNNs to be scaled beyond what is possible with DEER.
When selecting an approach, we offer the following advice: If rapid convergence is reliably observed, our experiments show that quasi-DEER provides the fastest convergence in terms of wall-clock time. However, if the dynamics are on the edge of stability, then ELK offers the most stable performance of these parallelized methods, but quasi-ELK could be faster in wall-clock time and just as stable. In such settings, it is worth sweeping the hyperparameter to choose the best version of ELK (note that for $\lambda=0$, ELK specializes to DEER). However, in the setting of chaotic dynamics, standard sequential evaluation may ultimately be faster.
Our experiments and these observations also highlight avenues for future research. While we found success with a diagonal approximation, structured approximations of the Jacobian that still admit fast parallelism but are more faithful approximations may allow for more accurate quasi-Newton steps to be taken. Secondly, quantifying the convergence rates of quasi-ELK would allow us to provide tighter bounds than those derived in Proposition 1. Finally, theoretically investigating whether further improvements to parallelized methods can prove faster than sequential evaluation for dynamical systems on the edge of stability, or whether there are fundamental limitations to the computational benefit of parallelization, are interesting questions for future work.
Acknowledgments
We thank John Duchi, David Zoltowski, and the members of the Linderman Lab for their helpful feedback. We thank Amber Hu for her work on preliminary versions of this project, and Leo Kozachkov for pointing out that the conditions we established for the merit function are equivalent to invexity. We thank the anonymous NeurIPS reviewers whose feedback improved this paper.
This work was supported by grants from the NIH BRAIN Initiative (U19NS113201, R01NS131987, & RF1MH133778), the NSF/NIH CRCNS Program (R01NS130789). X.G. would also like to acknowledge support from the Walter Byers Graduate Scholarship from the NCAA. S.W.L. is supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation. The authors have no competing interests to declare.
Some of the experiments were performed on the Sherlock cluster. We thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results.
@@ A.1 Proof of Proposition 1
Proposition 1 Undamped Newton’s method will converge to the true solution, $\mathbf{s}^{*}$, of the fixed-point equation(2) in at most $T$ Newton iterations, for any initial $\mathbf{s}^{(0)}$.
@@
Furthermore, we observe that if $\mathbf{s}_{t}^{(i)}=\mathbf{s}_{t}^{*}$ for all $t$ less than some $t^{(i)}$, then $\mathbf{r}_{t}(\mathbf{s}^{(i)})=\mathbf{0}$ for all $t<t^{(i)}$ by the definition of the residual in (1). Therefore, the DEER linear recurrence relation necessarily gives $\Delta\mathbf{s}^{(i+1)}_{t}=\mathbf{0}$ for all $t<t^{(i)}$. Furthermore, because $\mathbf{s}_{t^{(i)}}^{*}=f(\mathbf{s}_{t}^{(i)})$, it follows that $\Delta\mathbf{s}_{t^{(i)}}^{(i+1)}=-\mathbf{r}_{t^{(i)}}(\mathbf{s}^{(i)})=\mathbf{s}^{*}_{t^{(i)}}-\mathbf{s}^{(i)}_{t^{(i)}}$. Thus, it follows that after applying another Newton iteration that $\mathbf{s}_{t}^{(i+1)}=\mathbf{s}_{t}^{*}$ for all $t<t^{(i)}+1$. The global convergence result and bound on Newton iterates follows by induction. ∎
We note that this proof technique (induction) is very similar to that used to prove Proposition 1 of Shih et al. [50]. However, Shih et al. [50] proves a result about Picard iterations (zeroth order method). Our proof about the global convergence of Newton iterations contains the additional complication of dealing with a LDS (as a consequence of using a first order method). Note, that the global convergence comes from the zeroth order update; however, the first order gradient information can accelerate convergence. The fact that both Picard and Newton iterations have very similar proofs of global convergence suggests that they are closely related, and we will expand on this connection in future work.
Discussion of corollaries of Proposition 1
Corollaries of Proposition 1 include that the fixed-point iterations will converge (in at most $T$ iterations) to $\mathbf{s}^{*}$ even if the Jacobians $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$ are replaced by arbitrary matrices, and that we can reset the values of $\mathbf{s}^{(i)}_{t}$ for $t>i$ arbitrarily and still enjoy convergence.
In more detail, the elements in the sub-block-diagonal of $J:=\nicefrac{{\partial\mathbf{r}}}{{\partial\mathbf{s}}}$ can be replaced with arbitrary values – but the main block diagonal must remain as the identity and all other entries must be zero. Retaining convergence under modifications to the sub-block-diagonal portion is a corollary of Proposition 1, and can be seen from (6): If all the states up to and including position $t-1$ at the $(i)$th Newton iteration are correct, then the update in (6) at Newton iteration $(i+1)$ for position $t$ will use $\Delta\mathbf{s}_{t-1}^{(i+1)}=\mathbf{0}$ (no update is required at position $t-1$), and so the update to $\mathbf{s}_{t}^{(i+1)}$ no longer depends on the Jacobian.
We exploit this to develop quasi-DEER, retaining only the diagonal of the Jacobians. Doing so reduces the parallel scan from $\mathcal{O}(D^{3})$ work to $\mathcal{O}(D)$ work, making each iteration faster (while still admitting global convergence as above), but needs more Newton iterations to converge due to approximate updates. We find that this trade-off often yields a faster wall-clock time (cf. Figure 8).
Explicitly, the global convergence of quasi-DEER is a theoretical result (a corollary of Proposition 1), but the fast runtime of quasi-DEER in practice is an empirical result (cf. Figure 3).
A.2 The Merit Function Has No Local Minima or Saddle Points
@@
The merit function $\mathcal{L}(\mathbf{s})$ defined in (7) has a global minimum at the true trace $\mathbf{\mathbf{s}}^{*}$, satisfying $\mathcal{L}(\mathbf{s}^{*})=0$. It has no other critical points, i.e. no $\mathbf{s}$ such that $\nabla\mathcal{L}(\mathbf{s})=\mathbf{0}$ other than at the unique $\mathbf{s}^{*}$ for which $\mathbf{r}(\mathbf{s}^{*})=\mathbf{0}$.
@@
Since there are no critical points other than $\mathbf{s}^{*}$, the merit function cannot have local minima or saddle points. Since we have shown that every stationary point is a global minimum, it follows that the merit function $\mathcal{L}(\mathbf{s})$ is invex (cf. [74, 75]).
We also discuss further the uniqueness of the global minimizer $\mathbf{s}^{*}$ of the merit function $\mathcal{L}$.
For a deterministic forward function $f$ and fixed inputs there is a fixed sequence of states and outputs (note that any stochastic dynamics function can be reparameterized as deterministic by conditioning on the random inputs). Thus, $\mathbf{s^{*}}$ is the only sequence with zero residual (i.e. there is a unique sequence generated by the deterministic dynamics).
Furthermore, DEER cannot get stuck at any point that is not this sequence. We prove this in Proposition 1. Another way to see this however is that each update step (4) is equal to $\mathbf{J}^{-1}\mathbf{r}$. But, $\mathbf{J}$ is always invertible and so has trivial nullspace. Furthermore, the residual $\mathbf{r}$ can only be zero at the unique solution $\mathbf{s}^{*}$. Thus $\mathbf{J}^{-1}\mathbf{r}$ is nonzero everywhere except at the true solution, where it is zero. Thus, DEER cannot get stuck en route to finding the true and unique solution.
A.3 Kalman Filtering Damps the Eigenvalues of the Dynamics Matrices
A complementary perspective on how ELK results in more stable evaluation of nonlinear RNNs is to see how the Kalman filter damps the eigenvalues of the Jacobian matrices of the transition dynamics. We first provide a high-level overview, and then provide a more detailed derivation.
Overview
Let $\mathbf{A}_{t}$ be the Jacobians $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$ used in the linear recurrence relations and $\mathbf{b}_{t}$ be the offsets. Then the prediction step of the Kalman filter (ELK) is the same as DEER. However, after applying the update step in ELK (which imposes the trust region), we obtain a second linear recurrence relation where the linear operator is given by $\bm{\Gamma}_{t}\mathbf{A}_{T}$. Note that $\bm{\Gamma}_{t}$ is a symmetric positive definite matrix with eigenvalues bounded above by $\nicefrac{{1}}{{1+\lambda}}$. Thus, by the Spectral Theorem, it follows that the norms of the eigenvalues of $\bm{\Gamma}_{t}\mathbf{A}_{t}$ are bounded above by the max of the norms of the eigenvalues of $\mathbf{A}_{t}$, scaled by $\nicefrac{{1}}{{1+\lambda}}$. Note that larger $\lambda$ corresponds to more regularization/smaller trust region; and therefore correspondingly results in smaller effective eigenvalues in the scan. We recover DEER exactly if $\lambda=0$. Thus, while large eigenvalues in $A_{t}$ are the cause of the instability of DEER when evaluating unstable dynamical systems, ELK directly attenuates these large eigenvalues, explaining why the intermediate iterations using ELK remain stable.
Derivation
We define our dynamics used in Newton iteration $(i+1)$ as
$\mathbf{A}_{t}$ $=\dfrac{\partial f_{t+1}}{\partial\mathbf{s}}(\mathbf{s}_{t}^{(i)})$
$\mathbf{b}_{t}$ $=f_{t+1}(\mathbf{s}_{t}^{(i)})-\dfrac{\partial f_{t+1}}{\partial\mathbf{s}}(\mathbf{s}_{t}^{(i)})\mathbf{s}_{t}^{(i)}.$
Now $\mathbf{A}_{t}\in\mathbb{R}^{D\times D}$ and $\mathbf{b}_{t}\in\mathbb{R}^{D}$.
In line with considering the system as the LDS in (11), we set the process noise to be $\mathbf{I}_{D}$, and with the emissions governed by
$\mathbf{s}_{t}^{(i+1)}\sim\mathcal{N}(\mathbf{s}_{t}^{(i)},\sigma^{2}\mathbf{I}_{D}),$
where $\sigma^{2}$ controls the size of our trust region (note that in the notation of developed in Section 4.2 we have $\lambda=1/\sigma^{2}$).
In the notation of Murphy [76], we see that the predict step is
$\bm{\mu}_{(t+1)|t}$ $=\mathbf{J}_{t}\bm{\mu}_{t|t}+\mathbf{b}_{t}$
$\bm{\Sigma}_{(t+1)|t}$ $=\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+\mathbf{I}_{D}.$
Meanwhile, the update step is
$\bm{\mu}_{(t+1)|(t+1)}$ $=\bm{\mu}_{(t+1)|t}+\bm{\Sigma}_{(t+1)|t}(\bm{\Sigma}_{(t+1)|t}+\sigma^{2}\mathbf{I}_{D})^{-1}(\mathbf{y}_{t+1}-\bm{\mu}_{(t+1)|t})$
$\bm{\Sigma}_{(t+1)|(t+1)}$ $=\bm{\Sigma}_{(t+1)|t}-\bm{\Sigma}_{(t+1|t)}(\bm{\Sigma}_{(t+1|t)}+\sigma^{2}\mathbf{I}_{D})\bm{\Sigma}_{(t+1|t)}^{T}.$
To unpack this further, we first define the attenuation matrix
$\bm{\Gamma}_{t}=\sigma^{2}\Big{(}\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+(\sigma^{2}+1)\mathbf{I}_{D}\Big{)}^{-1}.$
Because $\bm{\Sigma}_{t|t}$ is a covariance matrix, it is also symmetric positive definite, and so $\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}$ is symmetric positive definite, and so all of its eigenvalues are greater than zero. Therefore, all the eigenvalues of $\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+(\sigma^{2}+1)\mathbf{I}_{D}$ are greater than $\sigma^{2}+1$.
We note that $\bm{\Gamma}_{t}$ is also symmetric and positive definite. Thus, by the Spectral Theorem, all eigenvalues of $\bm{\Gamma}_{t}$ are positive. By the above argument, the eigenvalues of $\bm{\Gamma}_{t}$ are all less than $\frac{\sigma^{2}}{1+\sigma^{2}}<1$.
Thus, we observe that the resulting filtering is given by the recurrence relation
$\bm{\mu}_{(t+1)|(t+1)}$ $=\overbrace{\bm{\Gamma}_{t}\mathbf{A}_{t}\bm{\mu}_{t|t}}^{\text{linear}}+\overbrace{\bm{\Gamma}_{t}\mathbf{b}_{t}+(\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+\mathbf{I}_{D})\Big{(}\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+(\sigma^{2}+1)\mathbf{I}_{D}\Big{)}^{-1}\mathbf{y}_{t+1}}^{\text{offset}}$
$\bm{\Sigma}_{(t+1)|(t+1)}$ $=\bm{\Gamma}_{t}(\mathbf{A}_{t}\bm{\Sigma}_{t|t}\mathbf{A}_{t}^{T}+\mathbf{I}_{D}).$
Given $\Big{\{}\bm{\Sigma}_{t|t}\Big{\}}_{t=0}^{T-1}$, we see that the filtered means (the updates for ELK) come from a linear recurrence relation with linear term $\bm{\Gamma}_{t}\mathbf{A}_{t}$.
We therefore compare the eigenvalues of $\bm{\Gamma}_{t}\mathbf{A}_{t}$ to eigenvalues of $\mathbf{A}_{t}$. Because $\bm{\Gamma}_{t}$ is symmetric positive definite, by the Spectral Theorem we can write $\bm{\Gamma}_{t}=\mathbf{Q}\bm{\Lambda}_{t}\mathbf{Q}^{T}$, where $\mathbf{Q}$ is an orthogonal (and therefore unitary) matrix, and $\bm{\Lambda}_{t}$ is a diagonal matrix where every entry is in $(0,1)$ (the entries of $\bm{\Lambda}_{t}$ are the eigenvalues of $\bm{\Gamma}_{t}$, which are greater than 00 by the Spectral Theorem and less than $\frac{\sigma^{2}}{1+\sigma^{2}}<1$ by the argument above).
Now, let’s consider any arbitrary unit vector $\mathbf{v}\in\mathbb{C}^{D}$, and let $\bm{\Lambda}_{t}^{\mathrm{max}}$ denote the maximum of the norms of all eigenvalues of $\mathbf{A}_{t}$. Then $\|\mathbf{A}_{t}\mathbf{v}\|_{2}\leq\bm{\Lambda}_{t}^{\mathrm{max}}$ by the definition of $\bm{\Lambda}_{t}^{\mathrm{max}}$. However, we want to know $\|\mathbf{Q}\bm{\Lambda}_{t}\mathbf{Q}^{T}\mathbf{A}_{t}\mathbf{v}\|_{2}$ for any arbitrary unit vector $\mathbf{v}\in\mathbb{R}^{D}$. However, we know that the action of a unitary matrix cannot change the 2-norm of a vector, so $\|\mathbf{Q}\bm{\Lambda}_{t}\mathbf{Q}^{T}\mathbf{A}_{t}\mathbf{v}\|_{2}=\|\bm{\Lambda}_{t}\mathbf{Q}^{T}\mathbf{A}_{t}\mathbf{v}\|_{2}$. Moreover, multiplying a vector by a diagonal matrix cannot increase the 2-norm of a vector by more than the absolute value of the diagonal matrix, which in the case of $\bm{\Lambda}_{t}$ is bounded above by $\nicefrac{{\sigma^{2}}}{{\sigma^{2}+1}}$. Thus, $\|\mathbf{Q}\bm{\Lambda}_{t}\mathbf{Q}^{T}\mathbf{A}_{t}\mathbf{v}\|_{2}\leq\frac{\sigma^{2}}{\sigma^{2}+1}\|\mathbf{A}_{t}\mathbf{v}\|$, or
$\|\mathbf{Q}\bm{\Lambda}_{t}\mathbf{Q}^{T}\mathbf{A}_{t}\mathbf{v}\|_{2}\leq\frac{\sigma^{2}}{1+\sigma^{2}}\bm{\Lambda}_{t}^{\mathrm{max}}$
for any unit vector $\mathbf{v}\in\mathbb{C}^{D}$. This highlights that we can interpret reducing $\sigma^{2}$ (reducing the size of the trust region and increasing stabilization) as directly attenuating the eigenvalues in the linear recurrence, helping to combat eigenvalues with large magnitude.
A.4 Scale-ELK
Motivated by our derivation in Appendix A.3, which shows that ELK reduces the magnitude of the eigenvalues of the Jacobian matrices in the transition dynamics, we recommend a more lightweight version of ELK which we call scale-ELK.
Scale-ELK uses a hyperparameter $k\in[0,1]$ (as opposed to $\lambda\in[0,\infty)$ used by ELK). Scale-ELK uses a linear dynamical system just like DEER, with the dynamics defined as
$\mathbf{A}_{t}$ $=(1-k)\dfrac{\partial f_{t+1}}{\partial\mathbf{s}}(\mathbf{s}_{t}^{(i)})$
$\mathbf{b}_{t}$ $=f_{t+1}(\mathbf{s}_{t}^{(i)})-(1-k)\dfrac{\partial f_{t+1}}{\partial\mathbf{s}}(\mathbf{s}_{t}^{(i)})\mathbf{s}_{t}^{(i)}.$
Thus, setting $k=0$ recovers DEER, while setting $k=1$ recovers a (computationally expensive form of) sequential evaluation. Ideally, $k$ is chosen to keep the magnitude of the eigenvalues of $\{A_{t}\}_{t=1}^{T}$ below 1. By Proposition 1, scale-ELK also enjoys global convergence.
Scale-ELK enjoys two primary benefits over ELK. First, an evaluation of scale-ELK uses fewer FLOPs than ELK, as scale-ELK is just parallelizing an LDS with ELK uses a parallelized Kalman filter. Second, the Kalman filter involves inverses which run the risk of introducing numerical instability, while scale-ELK avoids these complications.